OpenStudy (anonymous):

Linear Algebra: Let A be an n x n matrix. Show that if A^2 = 0, then I - A is nonsingular and (I - A)^-1 = I + A

OpenStudy (anonymous):

@jim_thompson5910 ?

OpenStudy (blockcolder):

You can show that (I-A)(I+A)=I from which you can conclude that (I+A)^(-1)=I-A and (I-A)^(-1)=I+A